The potential energy (in J) of a system in one dimension is given by: U(x) = 12 - 2x³ + 7x² - 4x. Which are the points of stable and unstable equilibrium for this potential function? Determine the work done in moving a particle in this potential from x = 1 m to x = 2 m
1
Expert's answer
2020-02-20T09:22:45-0500

 The work done in moving a particle in this potential from x = 1 m to x = 2 m:


"W=U(2)-U(1)"

"W=(12 - 2(2)^3 + 7(2)^2 - 4(2))-(12 - 2(1)^3 + 7(1)^2 - 4(1))=3\ J"

For the equilibrium:


"\frac{dU}{dx}=0=14x-6x^2-4"

The  points of equilibrium:


"x_1=\frac{1}{3}, x_2=2"

"\frac{d^2U}{dx^2}(x_1)=14-12\frac{1}{3}=10>0"

"\frac{d^2U}{dx^2}(x_2)=14-12(2)=-10<0"

x1 is the point of stable equilibrium, x2 is the point of unstable equilibrium.